Cardiarespiratry fitness is wicely recognized as a major camponent of overall physical well-being. Direct measurement consumiPlease explain steps and solve thanks
Answer #1

In a regression equation, the coefficient of a predictor can be interpreted in such a way that if predictor is Xi and its coefficient is ai , then holding all other variables constant, a 1 unit increase in Xi is going to cause ai units change in Y(response variable).

(a) Interpret  31,

Holding all other variables constant, a 1 kg increase in  weight will result in a 0.012 L/min increase in VO2max.

Interpret  \beta_3,

Holding all other variables constant, a 1 min increase in walk time will result in a 0.08 L/min decrease in VO2max.

(b) x1 = 79 , x2 = 24 , x3 = 12 , x4 = 134

Expected value of Y in this case is

y5.00.012 x 79 0.05 x 24 0.08 x 12 0.01 x 134

= 2.448

(c) When x1 = 79 , x2 = 24 , x3 = 12 , x4 = 134 , the Y is normally distributed

Y ~ N( lY = 2.448 , \sigma = 0.4)

We need to find,

P[ 1.928 < Y < 2.968 ]

= P[ (1.928 -2.448)/4 < (Y - 2.448)/4 < (2.968 - 2.448)/4]

= P[ -0.13 < Z < 0.13] ,{where Z = (Y - 2.448)/4 ~ N(0,1) }

= D(0.13) D(-0.13

= D(0.13) 1 (0.13) {since, Z is symmetrically distributed}

= 2D(0.13) 1

= 2 x 0.55172 - 1 {value is obtained from a standard normal table}

= 0.10344